OpenStudy (anonymous):

HELP WILL MEDAL AND FAN Which expression is always equivalent to m? m over three times six m over one times three m over three times three m over six times three

TheSmartOne (thesmartone):

Your answer is: m

OpenStudy (anonymous):

m over three times six m over one times three m over three times three m over six times three

TheSmartOne (thesmartone):

Are the answer choices supposed to be read like: for example A: \(\sf\Large \frac{m}{3}\times 6\) or \(\sf\Large \frac{m}{3 \times 6}\)

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

ok, so did they give you the answer choices in numbers or just all in words?

OpenStudy (anonymous):

numbers

TheSmartOne (thesmartone):

good

OpenStudy (anonymous):

okay

TheSmartOne (thesmartone):

do you know this: \(\sf\Large\frac{a}{b}\times c = \frac{a}{b}\times\frac{c}{1} = \frac{a\times c}{b \times 1}=\frac{a \times c}{b}\)

OpenStudy (anonymous):

nope

TheSmartOne (thesmartone):

you don't need to know it, but do you see how I multiplied the 2 fractions?

OpenStudy (anonymous):

yeah @TheSmartOne

TheSmartOne (thesmartone):

do you also know this then: \(\sf\Large\frac{a}{a}=1\)

OpenStudy (anonymous):

yep

TheSmartOne (thesmartone):

So what is your answer? Do the multiplication in each answer choice and simplify it. Which answer choice gives you back 'm' ?

OpenStudy (anonymous):

what does m stand or is what i dont get @TheSmartOne

TheSmartOne (thesmartone):

The m is in your answer choices and is in your question -.-

TheSmartOne (thesmartone):

it's the variable...

OpenStudy (anonymous):

no what is the number that m is replacing, i cant figure that out @TheSmartOne

TheSmartOne (thesmartone):

it does not replace any number

TheSmartOne (thesmartone):

https://www.mathsisfun.com/definitions/variable.html

OpenStudy (anonymous):

then how do you solve for m @TheSmartOne

TheSmartOne (thesmartone):

you are not solving for m.

TheSmartOne (thesmartone):

Read your own question. And no need to keep taggnig me. Each time you reply I'll get a notification and come back. You're not the only person I'm helping.

OpenStudy (anonymous):

then i dont get it

OpenStudy (anonymous):

please help me

TheSmartOne (thesmartone):

what don't you get?

OpenStudy (anonymous):

what to multiply

TheSmartOne (thesmartone):

\(\color{blue}{\text{Originally Posted by}}\) @TheSmartOne do you know this: \(\sf\Large\frac{a}{b}\times c = \frac{a}{b}\times\frac{c}{1} = \frac{a\times c}{b \times 1}=\frac{a \times c}{b}\) \(\color{blue}{\text{End of Quote}}\) I'll do A for you \(\sf\Large\frac{m}{3}\times 6 = \frac{m}{3}\times\frac{6}{1} = \frac{m\times 6}{3 \times 1}=\frac{m \times 6}{3}=2m\) \(\sf\Huge m \neq 2m\) So the answer is not A

OpenStudy (anonymous):

oh okay thaks

TheSmartOne (thesmartone):

np